# BOP-Distrib: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

### v3

@@ 2.5 YCB-V [48] Annotation Details
[…]
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 5.3$\times$5.3mm2. The pruning stage was not necessary for YCB-V, as objects are simpler, with much less occlusions.
@@ 6 Visualizing results by LiePose diffusion [19]
Figures 15, 16, 17 and 18 compare SpyroPose [9] and LiePose [19] results on objects with discrete and continuous symmetries. SpyroPose [9] rotations tends to be more precise than LiePose [19], but misses some of the modes. LiePose [19] tends to estimate continuous symmetries when the image produces discrete ones. These images are taken from a video compilation of SpyroPose [9] and LiePose [19] results also provided as supplementary material (BOP_Distrib_id8513_supp_distribution_comparison_SpyroPose_LiePose.mp4). Scenes with single instance of objects have been chosen, to facilitate visualization. We invite the reader to stop on some frames and check the differences in the estimates. Our ground truth distribution is displayed as the envelop for the rotation part and as red stars for the translation part.
@@ 7 Using Our Per-Image Annotations for Other Pose Estimation Datasets
We give here an illustration of the interest to reprocess their symmetries patterns. We take ITODD’s small validation set, for which the ground truth is public. Figure 19 presents our result for the star object new ground truth, as well as one case of the current best performer for Single Pose estimation (gpose2023 [49]) failing to align the holes. In the current version of BOP evaluation, this pose estimation is validated. With our annotations, it would be penalized.
@@
Acknowledgment
The authors thank Rasmus Laurvig Haugaard for providing SpyroPose [9] implementation for baseline evaluation.

### v4

@@ 2.5 YCB-V [54] Annotation Details
In our experiments, we sample surface points from the CAD models with a resolution of 1mm. For the pre-computation of elementary symmetries patterns for a given object, we use the per-object symmetries pattern given by the BOP challenge [21] as the initial symmetry candidates, with a tolerance factor $\epsilon\text{-sym}$ set to 2mm. The color tolerance is set in the HSV color space to have the chrominance on a single channel (hue) and only luminance on the other two (saturation and value). It is empirically set to 4 in hue and 0.1 in saturation and value.
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly $\times5.3$mm2. The pruning stage was not necessary for YCB-V, as objects are simpler, with much less occlusions.
@@ 6 Visualizing results by LiePose diffusion [23]
Figures 14, 15, 16 and 17 compare SpyroPose [13] and LiePose [23] results on objects with discrete and continuous symmetries. SpyroPose [13] rotations tends to be more precise than LiePose [23], but misses some of the modes. LiePose [23] tends to estimate continuous symmetries when the image produces discrete ones. These images are taken from a video compilation of SpyroPose [13] and LiePose [23] results also provided as supplementary material (BOP_Distrib_id8513_supp_distribution_comparison_SpyroPose_LiePose.mp4). Scenes with single instance of objects have been chosen, to facilitate visualization. We invite the reader to stop on some frames and check the differences in the estimates. Our ground truth distribution is displayed as the envelop for the rotation part and as red stars for the translation part.
@@
7 Metrics for evaluation: extended discussion
Rotation and translation errors are model-independent. [18], which led to BOP, states that ”fitness of object surface alignment is the main indicator of object pose quality, model-dependent pose error functions should be therefore preferred.” However, translation and rotation errors, as definer by [18]) with Equation 9 & Equation 10 can be exploited with our definitions of precision and recall over distributions (Equation 7 & Equation 8).
$\mathbf{d}_{\text{TE}}(\hat{\textbf{t}},\bar{\textbf{t}})=\|\bar{\textbf{t}}-\hat{\textbf{t}}\|_{2},$ (9)
$\mathbf{d}_{\text{RE}}(\hat{\textbf{R}},\bar{\textbf{R}})=\text{arccos}\left(\text{Tr}(\hat{\textbf{R}}\bar{\textbf{R}}-1)/2\right),$ (10)
where $\hat{\textbf{R}}$ and $\hat{\textbf{t}}$ are respectively the ground truth rotation and translation, and where $\bar{\textbf{R}}$ and $\bar{\textbf{t}}$ are respectively the estimated rotation and translation.
Table 4 is a reprocessing of pose distribution estimation method results, with precision and recall over distribution, as defined by Equation 7 & Equation 8, with translation and rotation errors from Equation 9 & Equation 10.
Methods $\mathbf{P_{RE}}$ $\mathbf{R_{RE}}$ $\mathbf{P_{TE}}$ $\mathbf{R_{TE}}$
SpyroPose [13] 73.2 69.1 43 66.9
LiePose [23] 68 91.1 46.2 92.9
Table 4: Precision/Recall over distribution for separate rotation and translation errors. We reprocess methods pose distribution estimates with decoupled rotation and translation errors.
8 Downstream Tasks Discussion
Our work implication on downstream task is two-fold. First, as highlighted by the BOP ranking changes in table 2, removing erroneous poses from the BOP ground truth implies a more reliable performance evaluation. Indeed, as illustrated by Fig. 11 of supp. mat., performance variations are related to poses that the current BOP ground-truth unduly classify as valid but are properly classified as invalid by our ground truth. This is due to overly lax ground truth annotation: some images are annotated as having multiple pose solutions due to the object symmetries whereas disambiguating elements breaking those symmetries are visible in the image. For downstream tasks, such as grasping or Augmented Reality, reliable ranking permits to choose the real best performer, and thus a higher success rate of the task.
The second implication on downstream tasks is related to Table 3, i.e. evaluating the ability of a method to determine the complete distribution of poses that explain the observed image. For applications such as grasping or Augmented Reality, if the object includes disambiguating elements, it implies that only one pose is valid for the task (e.g. a robot that should grab a mug at a specific position on the handle). However, if the observed image can be explained by multiple pose, it implies that disambiguating elements are not visible (e.g. the handle of the mug is not visible) and the downstream task is impossible to achieve from this unique viewpoint. A method that outputs the full distribution of poses provides the downstream task with the ability to determine if the task can be achieved (case of uni-modal distribution) or not (case of multi-modal solution). In such situation, a method that outputs a maximum of one pose does not permit to determine if the task can be achieved or not. Moreover, in such situation the complete distribution can help to determine the next best viewpoint to make the task feasible. Our distribution recall metric 8 evaluates the capacity of the estimation method to retrieve multi-modal distributions. This metric is a key indicator of pose distribution estimation methods for downstream tasks usability Figure 18 illustrates the robotic mug handle grabbing case.
Figure 18: Additional experiment for mug grasping task. We display multiple views of mug, associated to our pose distribution annotation. We are interested at pose estimation downstream tasks. We take the example of robotic grasping. We want the robot to grasp the mug by the handle. On view 1, the pose distribution is multi-modal, i.e. the object pose is ambiguous and the handle position in space cannot be accessed. For view 2, 3 and 4, the pose distribution is uni-modal, the image allows to estimate the pose of the mug without ambiguity. In such case, the downstream task of robotic handle grasping becomes feasible. Now, when it comes to evaluating pose distribution estimation methods against our ground truth, our recall metric evaluates the capacity of the estimation method to retrieve multi-modal distributions. This metric is a key indicator of pose distribution estimation methods for downstream tasks usability.
@@ 9 Using Our Per-Image Annotations for Other Pose Estimation Datasets
We give here an illustration of the interest to reprocess their symmetries patterns. We take ITODD’s small validation set, for which the ground truth is public. Figure 19 presents our result for the star object new ground truth, as well as one case of the current best performer for Single Pose estimation (gpose2023 [57]) failing to align the holes. In the current version of BOP evaluation, this pose estimation is validated. With our annotations, it would be penalized. New efforts at proposing challenging pose estimation datasets, such as [56], could be processed by our method. However, texture disambiguate a lot of the 3D models, and the scenes do not present much occlusion between objects.
@@
10 Discussion on Alignist [50]
Alignist [50] proposes to estimate rotation distribution for ambiguous object shapes from images. It was the first method that introduced a solution for supervising the training with a pseudo-ground truth generated rotation distribution. To do so, it resorts to a precomputation of such rotation distibution based on ground truth pose, rotation sampling, and SDF (Signed Distance Function) and Surfemb [12] features comparison. This precomputation is performed on renderings of single objects, and used to train a double MLP network to infer these distributions. The translation part of the pose is not considered. The test of the method on T-LESS [17] is conducted following Gilitschenski [11] protocol: it only processes single isolated objects on black background, and is evaluated with log likelihood.
In contrast, our annotation procedure does not rely on SurfEmb [12] comparisons which results are not guaranteed but it uses geometrical comparisons (see Equation 1). Moreover and unlike Alignist [50], our annotation procedure has a rejection mechanism for false visible points and false occluded points, as illustrated in Section 2.1 (see Section 3.3). This point is crucial to be able to generate a proper ground truth annotation. Finally, our approach does not need to retrain SurfEmb [12] to annotate a new dataset.
Alignist [50] and other pose distributions estimation methods would benefit from our more accurate pose distributions for their trainings. Finally, Alignist [50] method would benefit from our evaluation framework (see Section 5.4). For that though, Alignist [50] would need to be tested on the full T-LESS [17] test set (and not just Gilitschenski [11] protocol that excludes external occlusions). We could not conduct such tests as the codes and results are not public.